OpenStudy (anonymous):

write the equation of a line that passes through the points (-1,7)(-2,5)

OpenStudy (anonymous):

I recommend using point-slope or slope-intercept form. Point slope: (y1-y2)=m(x1-x2) Slope Intercept: y=mx+b m represents the slope of the equation, and b represents y intercept. The numbers for y and x are to imply the coordinates of each ordered pair. For instance: -1 could be x1 and -2 can be x2, or the other way around. The important thing though, is that if you label -1 as x1, then 7 MUST be labeled y1. The same reasoning goes for -2 being x2. 5 MUST be labeled y2.

OpenStudy (luigi0210):

Slope formula: \(\Large \frac{y_{2}-y_{1}}{x_{2}-x{1}}\) Then the point-slope formula \(\Large y-y_{1}=m(x-x_{1})\)